Which causes a largest change in pH on addition to 20 ml 0.1 M acetic acid solution? 


        (A)    Addition of distilled water


        (B)    Addition of 20 mol of 0.01 MCH3COOH


        (C)    Addition of 20 ml of 0.1 M HCl


        (D)    Addition of 20 ml of 0.1 M NaOH


Grade:10

1 Answers

Navjyot Kalra
10 years ago

Moles of AcOH = 20 × 10−3 × 0.1 = 2 ×10−2

 

        Moles of NaOH = 20 × 10−3 × 0.1 = 2 × 10−3

 

        pH will change from acidic side to basic